# Vision-Language Alignment from Compressed Image Representations using 2D Gaussian Splatting

### v1

@@ Abstract
Figure 1: (a) Architecture of an autoregressive visual language model (VLM) (b) 2DGS adaptation of contrastive language-image pre-training (CLIP) (c) Illustration of 2D Gaussian Splatting (2DGS) for image fitting. Each image is represented as a sparse mixture of anisotropic Gaussians parameterized by position, covariance, and color. Summing contributions from all splats reconstructs the original image (with a minor and configurable degradation loss), enabling compact, spatially adaptive representations.
@@ 1 Introduction
Expressive yet compact visual representations are central to progress in computer vision and multimodal learning. Recent advances in large multimodal models (LMMs), particularly visual language models (VLMs), have extended the reasoning capabilities of large language models beyond text, enabling impressive capabilities in visual question answering, image captioning, and cross-modal understanding. VLMs leverage a shared embedding space to enable joint reasoning over text, images, and other modalities [19, 15, 16, 18, 31, 20], as illustrated by Fig. 1(a). Central to this paradigm are vision encoders that map raw images into sequences of visual tokens, semantically aligned with textual embeddings. The canonical pipeline starts from (often high-resolution) RGB images, which are patchified using a uniform grid and routinely projected into hundreds or thousands of tokens. These encoders typically adopt a vision transformer (ViT) backbone, trained via contrastive language–image pre-training (CLIP) on large image–caption datasets [26].
While effective, this paradigm suffers from two systemic inefficiencies. (1) Reliance on high-density RGB inputs imposes subtantial data-transfer and energy costs in edge–cloud deployments: even when compressed into formats such as JPEG for transfer, visual data remains bandwidth-intensive and requires costly decoding before processing. (2) Patch-based tokenization causes a severe token explosion: whereas a text prompt may require only a few dozen tokens, vision encoders routinely emit hundreds or thousands per image, many of which are redundant. This redundancy inflates memory and compute costs due to the quadratic scaling of attention, threatening the efficiency and sustainability of large-scale multimodal training and deployment.
We hypothesize that part of this inefficiency stems from a mismatch between human-centric visual formats and the needs of machine learning systems. RGB pixel arrays are tailored for human perception, with dense spatial correlations, but may include far more detail than models require. Unlike humans, who rely on full-resolution imagery, learning systems could benefit from more abstract representations that emphasize semantic and structural content while discarding perceptually oriented redundancy. This perspective raises a central research question: instead of transmitting perceptually complete images intended for human viewing, can we design compact, semantically rich representations that preserve information critical for downstream learning while substantially reducing transmission overhead?
@@
  • Large-Scale CLIP Adaptation for Gaussian Splat Inputs: We systematically adapt the CLIP framework to operate on 2DGS representations, via a novel GS stem, including Fourier positional encodings, log-scaled covariance representations, and a Perceiver Resampler pre-processing unit on GS inputs. Our approach leverages frozen RGB-based transformer backbones while introducing lightweight, splat-aware components, enabling efficient transfer learning from established vision encoders.
Our experimental evaluation on a 12.8M subset of DataComp [8] provides the first systematic study of GS-based vision encoders in zero-shot classification tasks, a widely used proxy for assessing vision–language alignment. While GS-based encoders currently trail RGB-based models in zero-shot accuracy, they demonstrate clear viability as compressed, transferable representations. Unlike RGB pipelines, which have benefited from years of architectural refinements and mature training recipes, 2DGS-based CLIP pipelines are in their early stages and poised for rapid improvement. We position this work not as an incremental patch-based encoder improvement, but as a first step toward a fundamentally different paradigm: building multimodal systems from inherently compact, adaptive representations rather than post-hoc compression of tokens.
@@ 2 Motivation and Related Work
The Edge-Cloud Data Transfer Challenge Modern multimodal pipelines increasingly depend on edge–cloud frameworks, where bandwidth-limited devices capture high-resolution visual data and transmit it to cloud-based vision encoders for semantic processing. Yet, this paradigm hides a critical inefficiency: transmitting dense RGB-based visual data remains energy-intensive, particularly for resource-constrained edge devices. Empirical studies indicate that mobile uplink transmission (the primary mode for edge devices) alone consumes approximately 0.1–0.2 kWh per GB, (as compared to $\sim$0.03 kWh per GB for fixed broadband connections) [1]. At scale, this may translate to tens of kilowatt-hours for just a few gigabytes of visual data. For instance, a single hour of 1080p video transmission consumes 0.38-0.68 kWh over mobile networks, energetically equivalent to powering a typical home for hours [9]. As such, the energy and infrastructure costs of naive RGB data transfer can quickly dominate operational budgets in real-world multimodal systems.
@@
The Token Explosion Problem and the Increasing Strain on Context Lengths The second challenge emerges in downstream processing, where current vision encoders generate sequences of hundreds or thousands of tokens per image. For example, LLaVA-1.5 [19] emits 576 tokens for a $336\times 336$ image, while LLaVA-NeXT [20] outputs over 2,880 tokens for $672\times 672$ inputs, compared to about 20–50 tokens for a typical sentence in a text prompt. This visual information density creates severe computational bottlenecks, as self-attention scales quadratically with sequence length. To address this, recent work has focused on post-hoc token reduction. PruMerge [28] and VisionZip [33] implement text-agnostic selection strategies, with the former requiring LLM-level fine-tuning and the latter projector-level fine-tuning for optimal performance. ToMe [3] integrates token merging directly into encoder blocks via cosine-similarity–based selection. Progressive sparsification methods such as FastV [4] and SparseVLM [36] prune tokens layer by layer, guided by attention weights or cross-modal relevance. More recently, cluster-based aggregation [25] has shown that simple, finetuning-free grouping can outperform prior methods, with minimal loss in downstream performance. Strikingly, discarding up to 89% of visual tokens often has little effect on VQA benchmarks [17, 12, 29, 22, 21, 7, 34, 10], and in some cases even improves accuracy. Even naive strategies like spatial or random sampling perform competitively, highlighting the vast redundancy of current visual tokens.
While effective in reducing sequence length, these approaches remain constrained by patch-based tokenization; they are, in essence, band-aid solutions to an architectural inefficiency. The fact that most tokens can be discarded without harming performance underscores a deeper issue: today’s encoders may be producing large volumes of redundant, semantically shallow information. This not only inflates compute, but also raises a broader question: are patch-based encoders truly the right substrate for multimodal alignment? These findings suggest an opportunity to move beyond post-hoc pruning toward representations that are inherently compact, semantically expressive, and computationally efficient.
Compressed and Alternative Image Representations The limitation of post-hoc token reduction extends beyond compute efficiency. Such methods assume that optimal visual representations must originate from dense pixel grids, and only later compress them. This paradigm overlooks the possibility that alternative input representations could be inherently more compact, expressive, and better aligned with the statistical patterns neural networks exploit.
@@
Traditional codecs such as JPEG achieve high compression ratios but yield representations that lack interpretability for learning systems, though some work has explored making compressed formats more machine-friendly. Implicit neural representations (INRs) instead encode signals as continuous functions parameterized by neural networks, achieving high fidelity in compression and reconstruction [30]. Recent studies have even treated INR parameters as input embeddings for downstream tasks, with promising results [2, 6]. However, INRs remain computationally costly to optimize and are ill-suited for large-scale training pipelines.
A recent alternative is 2D Gaussian Splatting (2DGS): a 2D adaptation [35] of 3D Gaussian Splatting for real-time radiance field rendering [14]. 2DGS parameterizes images as sparse mixtures of anisotropic Gaussians, each defined by spatial position, covariance, and color. Unlike uniform patch grids, 2DGS adapts its representational density to local image complexity. Early studies show state-of-the-art image fitting with orders-of-magnitude faster encoding than INRs, making it far more practical for large-scale use. Preliminary work also suggests its viability as both a codec and a feature representation for small-scale classification tasks.
Yet, the potential of 2DGS as a learnable substrate for large-scale multimodal alignment remains unexplored. Meeting this challenge requires representations that capture not only visual fidelity but also semantic alignment with text. We argue that the compactness and adaptivity of 2DGS make it a particularly promising candidate for efficient, semantically meaningful visual embeddings in multimodal systems.
3 Preliminaries
3.1 2D Gaussian Splatting as a Compressor
@@ 3 Preliminaries
Building on 3D Gaussian Splatting (3DGS) for real-time novel-view synthesis [14], [35] introduced GaussianImage, adapting the concept for single-image compression. Instead of heavy 3D Gaussians requiring tens of parameters, GaussianImage uses compact 2D Gaussians parameterized by only eight values: 2D position, anisotropic covariance, and color, as illustrated by Fig. 1(c). Images are reconstructed by summing contributions from all Gaussians, following Equation 1.
@@
$\begin{split}\hat{I}\!\left(x,y;\,\{\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i},\mathbf{c}_{i}\}_{i=1}^{n}\right)&=\sum_{i=1}^{n}\mathbf{c}_{i}\,\exp\!\Bigg\{-\tfrac{1}{2}\left(\begin{bmatrix}x\\ y\end{bmatrix}-\boldsymbol{\mu}_{i}\right)^{\!\top}\\ &\qquad\boldsymbol{\Sigma}_{i}^{-1}\left(\begin{bmatrix}x\\ y\end{bmatrix}-\boldsymbol{\mu}_{i}\right)\Bigg\},\end{split}$ (1)
@@
2D Gaussian Splatting eliminates the need for depth sorting, enabling extremely fast rendering (1,5001{,}5002,0002{,}000 FPS, or equivalently over 3X the decoding speed of JPEG), substantial memory savings, and $\sim 5\times$ faster training compared to implicit neural representations (INRs) such as WIRE [27] or I-NGP [24] [35].
@@
For further compression, GaussianImage employs a quantization pipeline (16-bit floats for positions, 6-bit integer covariance quantization, residual vector quantization for color, and partial bits-back coding). This achieves rate–distortion performance comparable to INR-based codecs while offering orders-of-magnitude faster decoding [35]. Following this, [37] proposed LIG (Large Images are Gaussians), a hierarchical two-level scheme that first fits low-frequency content and then refines high-frequency details, significantly improving scalability to ultra-high-resolution images. Extending to video, GSVC [32] predicts Gaussians sequentially across frames, prunes low-impact ones, adds new splats as needed, and reinitializes at keyframes, achieving AV1/VVC-level rate–distortion performance while decoding 1080p video at  1,500 FPS [32].
3.2 2D Gaussian Splatting as a Learnable Representation
@@
Beyond compression, 2DGS has been explored as a learnable substrate for vision tasks. GViT [11] replaces pixel patches with hundreds of Gaussians, and jointly optimizes them with a Vision Transformer. A differentiable renderer enforces image reconstruction, while classification gradients push Gaussians toward discriminative regions, yielding a compact, interpretable representation that matches ViT-B performance on ImageNet-1K (76.9% top-1) supervised classification. [5] proposes GaussianToken, which encodes images as Gaussians with continuous spatial parameters and feature coefficients, quantized via a VQ codebook and concatenated with spatial attributes. This hybrid token design improves reconstruction fidelity over VQ-VAE tokenizers.
Yet, existing efforts remain limited to small-scale settings. The role of 2DGS as a foundation for large-scale multimodal alignment, where representations must balance compactness, semantic richness, and learnability, has not been systematically investigated. Our work addresses this gap, providing the first large-scale study of 2DGS within vision–language pre-training.
In the context of our work, the key takeaway is that 2DGS provides a compact, spatially adaptive, and computationally efficient image representation. Rather than using 2DGS solely as a codec, we investigate its role as a learnable substrate for multimodal alignment, asking whether its promising properties can support large-scale vision–language pre-training.
3.3 Contrastive Language-Image Pre-training (CLIP)
Vision–language alignment is commonly achieved through Contrastive Language–Image Pre-Training (CLIP), which trains a dual-encoder architecture on image–caption pairs. A text encoder and a vision encoder are optimized jointly so that embeddings of aligned pairs are pulled closer, and mismatched pairs pushed apart under a contrastive loss [26]. CLIP has become the standard paradigm, demonstrating strong performance across a wide range of downstream multimodal tasks. The CLIP framework spans a family of models from lightweight 25M-parameter encoders to larger 300M+ parameter variants such as ViT-L/14, all of which require hundreds of millions of training samples and significant compute to reach high-quality alignment.
In this work, we retain the CLIP training recipe but investigate whether 2D Gaussian Splatting can serve as an alternative visual substrate. Using the OpenCLIP implementation [23], we directly compare RGB- and 2DGS-based encoders under identical training conditions, isolating the effect of the representation itself. Our approach further reuses frozen RGB-pretrained backbones with lightweight splat-aware input modules, enabling efficient adaptation while substantially reducing trainable parameters.
@@ 4.1 Challenges
Scaling 2D Gaussian Splatting (2DGS) for vision–language alignment faces three main challenges. First, without efficient batching, fitting a single image requires multiple forward and backward passes through a differentiable renderer, taking several seconds and making large-scale training impractical. Second, the current ecosystem lacks maturity: open implementations such as GaussianImage lack batch optimization and exhibit poor GPU utilization, further limiting throughput. Finally, achieving useful compression requires careful trade-offs: reducing the number of Gaussians lowers cost, but naïve pruning can harm perceptual quality and discard semantically important information. Together, these issues restrict the applicability of 2DGS in multimodal pipelines and motivate our system- and algorithm-level optimizations.
@@ 4.2 Specialized CUDA Kernels for Scalable 2DGS
Our implementation builds on the open-source GaussianImage codebase [35], which first introduced 2D Gaussian Splatting for high-speed image representation. Subsequent work by [37] streamlined this design, restructuring the kernels and data flow specifically for 2D operations and thereby yielding a cleaner, more efficient foundation for image representation.
This codebase is structured as follows: the forward pass is decomposed into two phases: (1) a projection stage, wherein each Gaussian primitive is projected onto a 2D grid of pixels. The grid is tiled across thread blocks, where each block corresponds to a 2D tile of the output image. Within a block, individual threads are assigned to pixels, computing Gaussian weights and contributions in parallel. This design ensures that Gaussian evaluation is parallelized across both the spatial dimension (pixels) and across different Gaussians, and (2) a rasterization stage, wherein the projected Gaussians are accumulated to produce the final output image at each iteration. This stage handles blending and weighting of Gaussians for each pixel. Unlike projection, where computation is Gaussian-centric, rasterization is pixel-centric, with threads accumulating contributions into shared output buffers. The backward pass is handled by a dedicated kernel that computes gradients of the loss with respect to Gaussian parameters (positions, covariances, colors). Workload partitioning is inverted relative to the forward pass: each block is assigned to a 2D tile of the input grid, and threads within the block compute pixel-level derivatives. This tiling strategy balances workload across warps and ensures efficient coalesced memory access when accumulating gradients.
Figure 2: Speedup results achieved by our CUDA kernels compared to the [37] baseline, following our batch-aware implementation. Speedups are presented for various batch sizes and Gaussian counts for image resolutions of 224x224. For a batch size of 4096 and 400 Gaussian points per image, we observe a 90.3X speedup compared to the baseline.
@@
With these modifications, the codebase now supports high-speed processing of thousands of inputs concurrently, making the implementation suitable for large-scale training pipelines. Notably, our enhanced codebase enables up to 90X speedup in fitting time compared to the baseline open-sourced codebase by [37] with  97% GPU utilization, profiled using NVIDIA Nsight Compute. Fig. 2 presents speedup results for various batch sizes and gaussian counts. Our optimized 2DGS pipeline, including both our algorithmic and system-level optimizations will soon be open-sourced to enable further explorations of 2DGS.

### v2

# GaussianVision: Vision-Language Alignment from Compressed Image Representations using 2D Gaussian Splatting

@@ Abstract
Figure 1: (a) Illustration of 2D Gaussian Splatting (2DGS) for image fitting. Each image is represented as a sparse mixture of anisotropic Gaussians parameterized by position, covariance, and color. Summing contributions from all splats reconstructs the original image (with a minor and configurable degradation loss), enabling compact, spatially adaptive representations. (b) 2DGS adaptation of contrastive language-image pre-training (CLIP). (c) Architecture of an autoregressive visual language model (VLM). (d) Architecture of our 2DGS-adapted CLIP pipeline: a splat-aware stem embeds a configurable number of Gaussian points using Fourier features, log scaling, normalization layers, and projections. These embeddings are processed by a majoritarily frozen RGB pretrained transformer.
@@ 1 Introduction
Expressive yet compact visual representations are central to progress in computer vision and multimodal learning. Recent advances in large multimodal models (LMMs), particularly visual language models (VLMs), have extended the reasoning capabilities of large language models beyond text, enabling impressive capabilities in visual question answering, image captioning, and cross-modal understanding. VLMs leverage a shared embedding space to enable joint reasoning over text, images, and other modalities [Liu2023ImprovedBaselines, Li2024LLaVAOneVision, Li2023BLIP2, Lin2024VILA, AdeptAI2023Adept, Liu2024LLaVANext], as illustrated by Fig. 1c. Central to this paradigm are vision encoders that map raw images into sequences of visual tokens, semantically aligned with textual embeddings. The canonical pipeline starts from (often high-resolution) RGB images, which are patchified using a uniform grid and routinely projected into hundreds or thousands of tokens. These encoders typically adopt a vision transformer (ViT) backbone, trained via contrastive language–image pre-training (CLIP) on large image–caption datasets [Radford2021CLIP]. While effective, this paradigm suffers from two systemic inefficiencies. (1) Reliance on high-density RGB inputs imposes subtantial data-transfer and energy costs in edge–cloud deployments: even when compressed into formats such as JPEG for transfer, visual data remains bandwidth-intensive and requires costly decoding before processing. (2) Patch-based tokenization causes a severe token explosion: whereas a text prompt may require only a few dozen tokens, vision encoders routinely emit hundreds or thousands per image, many of which are redundant. This redundancy inflates memory and compute costs due to the quadratic scaling of attention, threatening the efficiency and sustainability of large-scale multimodal training and deployment. We hypothesize that part of this inefficiency stems from a mismatch between human-centric visual formats and the needs of machine learning systems. RGB pixel arrays are tailored for human perception, with dense spatial correlations, but may include far more detail than models require. Unlike humans, who rely on full-resolution imagery, learning systems could benefit from more abstract representations that emphasize semantic and structural content while discarding perceptually oriented redundancy. This perspective raises a central research question: instead of transmitting perceptually complete images intended for human viewing, can we design compact, semantically rich representations that preserve information critical for downstream learning while substantially reducing transmission overhead?
@@
  • Large-Scale CLIP Adaptation for Gaussian Splat Inputs: We systematically adapt the CLIP framework to operate on 2DGS representations, via a novel GS stem and a lightweight two-stage framework for efficient transfer learning from established vision encoders.
We position this work not as an incremental patch-based encoder improvement, but as a first step toward a fundamentally different paradigm: building multimodal systems from inherently compact, adaptive representations rather than post-hoc compression of tokens.
@@ 2 Motivation and Related Work
The Edge-Cloud Data Transfer Challenge Modern multimodal pipelines increasingly depend on edge–cloud frameworks, where bandwidth-limited devices capture high-resolution visual data and transmit it to cloud-based vision encoders for semantic processing. Yet, this paradigm hides a critical inefficiency: transmitting dense RGB-based visual data remains energy-intensive, particularly for resource-constrained edge devices. Empirical studies indicate that mobile uplink transmission (the primary mode for edge devices) alone consumes approximately 0.1–0.2 kWh per GB, (as compared to $\sim$0.03 kWh per GB for fixed broadband connections) [EU_ICT_Impact_2020]. At scale, this may translate to tens of kilowatt-hours for just a few gigabytes of visual data. For instance, a single hour of 1080p video transmission consumes 0.38-0.68 kWh over mobile networks, energetically equivalent to powering a typical home for hours [GSMA_2025_mobile_data_energy]. As such, the energy and infrastructure costs of naive RGB data transfer can quickly dominate operational budgets in real-world multimodal systems.
@@
  The Token Explosion Problem and the Increasing Strain on Context Lengths The second challenge emerges in downstream processing, where current vision encoders generate sequences of hundreds or thousands of tokens per image. For example, LLaVA-1.5 [Liu2023ImprovedBaselines] emits 576 tokens for a $336\times 336$ image, while LLaVA-NeXT [Liu2024LLaVANext] outputs over 2,880 tokens for $672\times 672$ inputs, compared to about 20–50 tokens for a typical sentence in a text prompt. This visual information density creates severe computational bottlenecks, as self-attention scales quadratically with sequence length. To address this, recent work has focused on post-hoc token reduction. PruMerge [Shang2024prumerge] and VisionZip [Yang2024visionzip] implement text-agnostic selection strategies, with the former requiring LLM-level fine-tuning and the latter projector-level fine-tuning for optimal performance. ToMe [tome] integrates token merging directly into encoder blocks via cosine-similarity–based selection. Progressive sparsification methods such as FastV [Chen2024fastv] and SparseVLM [Zhang2025sparsevlm] prune tokens layer by layer, guided by attention weights or cross-modal relevance. More recently, cluster-based aggregation [Omri2025tokens] has shown that simple, finetuning-free grouping can outperform prior methods, with minimal loss in downstream performance. Strikingly, discarding up to 89% of visual tokens often has little effect on VQA benchmarks [Li2023POPE, hudson2019gqa, singh2019textvqa, Lu2022ScienceQA, Liu2024MMBench, Fu2023MME, Yu2024MMVet, gurari2018vizwiz], and in some cases even improves accuracy. Even naive strategies like spatial or random sampling perform competitively, highlighting the vast redundancy of current visual tokens. While effective in reducing sequence length, these approaches remain constrained by patch-based tokenization; they are, in essence, band-aid solutions to an architectural inefficiency. The fact that most tokens can be discarded without harming performance underscores a deeper issue: today’s encoders may be producing large volumes of redundant, semantically shallow information. This not only inflates compute, but also raises a broader question: are patch-based encoders truly the right substrate for multimodal alignment? These findings suggest an opportunity to move beyond post-hoc pruning toward representations that are inherently compact, semantically expressive, and computationally efficient.
@@
  Compressed and Alternative Image Representations The limitation of post-hoc token reduction extends beyond compute efficiency. Such methods assume that optimal visual representations must originate from dense pixel grids, and only later compress them. This paradigm overlooks the possibility that alternative input representations could be inherently more compact, expressive, and better aligned with the statistical patterns neural networks exploit. Traditional codecs such as JPEG achieve high compression ratios but yield representations that lack interpretability for learning systems, though some work has explored making compressed formats more machine-friendly. Implicit neural representations (INRs) instead encode signals as continuous functions parameterized by neural networks, achieving high fidelity in compression and reconstruction [sitzmann2020SIREN]. Recent studies have even treated INR parameters as input embeddings for downstream tasks, with promising results [Bauer2023SpatialFuncta, Dupont2022FromDataToFuncta]. However, INRs remain computationally costly to optimize and are ill-suited for large-scale training pipelines. A recent alternative is 2D Gaussian Splatting (2DGS): a 2D adaptation [Zhang2024GaussianImage] of 3D Gaussian Splatting for real-time radiance field rendering [Kerbl2023ThreeDGaussianSplatting].
3 2D Gaussian Splatting Preliminaries
@@
Building on 3D Gaussian Splatting (3DGS) for real-time novel-view synthesis [Kerbl2023ThreeDGaussianSplatting], [Zhang2024GaussianImage] introduced GaussianImage, adapting the concept for single-image compression. Instead of heavy 3D Gaussians requiring tens of parameters, GaussianImage uses compact 2D Gaussians parameterized by only eight values: 2D position, anisotropic covariance, and color, as illustrated by Fig. 1a. Images are reconstructed by summing contributions from all Gaussians, following Equation 1.
@@
$\begin{split}\hat{I}\!\left(x,y;\,\{\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i},\mathbf{c}_{i}\}_{i=1}^{n}\right)&=\sum_{i=1}^{n}\mathbf{c}_{i}\,\exp\!\Bigg\{-\tfrac{1}{2}\left(\begin{bmatrix}x\\ y\end{bmatrix}-\boldsymbol{\mu}_{i}\right)^{\!\top}\\ &\hskip 18.49988pt\boldsymbol{\Sigma}_{i}^{-1}\left(\begin{bmatrix}x\\ y\end{bmatrix}-\boldsymbol{\mu}_{i}\right)\Bigg\},\end{split}$ (1)
@@
2D Gaussian Splatting eliminates the need for depth sorting, enabling extremely fast rendering (1,5001{,}5002,0002{,}000 FPS, or equivalently over 3X the decoding speed of JPEG), substantial memory savings, and $\sim 5\times$ faster training compared to implicit neural representations (INRs) such as WIRE [saragadam2023wirewaveletimplicitneural] or I-NGP [M_ller_2022] [Zhang2024GaussianImage].
@@
For further compression, GaussianImage employs a quantization pipeline (16-bit floats for positions, 6-bit integer covariance quantization, residual vector quantization for color, and partial bits-back coding). This achieves rate–distortion performance comparable to INR-based codecs while offering orders-of-magnitude faster decoding [Zhang2024GaussianImage]. Following this, [Zhu2025LIG] proposed LIG (Large Images are Gaussians), a hierarchical two-level scheme that first fits low-frequency content and then refines high-frequency details, significantly improving scalability to ultra-high-resolution images. Extending to video, GSVC [Wang2025GSVC] predicts Gaussians sequentially across frames, prunes low-impact ones, adds new splats as needed, and reinitializes at keyframes, achieving AV1/VVC-level rate–distortion performance while decoding 1080p video at  1,500 FPS.
@@
Beyond compression, 2DGS has been explored as a learnable substrate for vision tasks. GViT [Hernandez2025GViT] replaces pixel patches with hundreds of Gaussians, and jointly optimizes them with a Vision Transformer. A differentiable renderer enforces image reconstruction, while classification gradients push Gaussians toward discriminative regions, yielding a compact, interpretable representation that matches ViT-B performance on ImageNet-1K (76.9% top-1) supervised classification. [Dong2025GaussianToken] proposes GaussianToken, which encodes images as Gaussians with continuous spatial parameters and feature coefficients, quantized via a VQ codebook and concatenated with spatial attributes. This hybrid token design improves reconstruction fidelity over VQ-VAE tokenizers. Yet, existing efforts remain limited to small-scale settings. Our work addresses this gap, providing the first large-scale study of 2DGS within vision–language pre-training, asking whether its promising properties can support vision–language alignment at scale.
@@ 4.1 Challenges
Large-scale vision–language alignment requires fitting millions of images, but existing 2DGS pipelines are still too slow, leaving substantial headroom in batching, memory layout, and GPU utilization. As a result, naïvely applying 2DGS at scale is infeasible without both algorithmic refinements and systems-level restructuring. We therefore begin our study by developing a set of optimizations that make high-throughput 2DGS fitting practical at the scale required for downstream alignment.
@@ 4.2 Specialized CUDA Kernels for Scalable 2DGS
Our implementation builds on the open-source GaussianImage codebase [Zhang2024GaussianImage], which first introduced 2D Gaussian Splatting for high-speed image representation. Subsequent work by [Zhu2025LIG] streamlined this design, restructuring the kernels and data flow specifically for 2D operations and thereby yielding a cleaner, more efficient foundation for image representation. This codebase is structured as follows: the forward pass is decomposed into two phases: (1) a projection stage, wherein each Gaussian primitive is projected onto a 2D grid of pixels. The grid is tiled across thread blocks, where each block corresponds to a 2D tile of the output image. Within a block, individual threads are assigned to pixels, computing Gaussian weights and contributions in parallel. This design ensures that Gaussian evaluation is parallelized across both the spatial dimension (pixels) and across different Gaussians, and (2) a rasterization stage, wherein the projected Gaussians are accumulated to produce the final output image at each iteration. This stage handles blending and weighting of Gaussians for each pixel. Unlike projection, where computation is Gaussian-centric, rasterization is pixel-centric, with threads accumulating contributions into shared output buffers. The backward pass is handled by a dedicated kernel that computes gradients of the loss with respect to Gaussian parameters (positions, covariances, colors). Workload partitioning is inverted relative to the forward pass: each block is assigned to a 2D tile of the input grid, and threads within the block compute pixel-level derivatives. This tiling strategy balances workload across warps and ensures efficient coalesced memory access when accumulating gradients.
Figure 2: Speedup results achieved by our CUDA kernels compared to the [Zhu2025LIG] baseline, following our batch-aware implementation. Speedups are presented for various batch sizes and Gaussian counts for image resolutions of 224x224. For a batch size of 4096 and 400 Gaussian points per image, we observe a 90.3X speedup compared to the baseline.
@@
With these modifications, the codebase now supports high-speed processing of thousands of inputs concurrently, making the implementation suitable for large-scale training pipelines. Notably, our enhanced codebase enables up to 90X speedup in fitting time compared to the baseline open-sourced codebase by [Zhu2025LIG] with  97% GPU utilization, profiled using NVIDIA Nsight Compute. We provide additional details on our CUDA optimizations in Appendix C. Fig. 2 presents speedup results for various batch sizes and gaussian counts. Our optimized 2DGS pipeline, including both our algorithmic and system-level optimizations will soon be open-sourced.